# TEXasGAN: Tactile Texture Exploration and Synthesis System Using Generative Adversarial Network

### v2

@@ V-A Encoded Latent Space
V-B Problems in Other Alternatives
In addition to the network model used in TEXasGAN, we implemented various other generators or decoders for comparison, including basic GAN, autoencoder; Variational Autoencoder (VAE) to learn the distribution; AAE to align the distribution, Conditional GAN (CGAN), and ACGAN to incorporate conditional information. However, these models faced challenges operating within our optimization framework, making direct comparisons unfeasible here.
Initially, we experimented with models based on the autoencoder architecture, including autoencoders, VAEs, and AAEs. While these models were relatively easy to train, they suffered from severe over-smoothing issues. Given our need for feature vectors suitable for optimization, we targeted a relatively small feature space dimension, such as the 128 dimensions used in this study. However, this dimensionality proved inadequate for the original data, resulting in significant information loss when encoded into this latent space. Consequently, the generated spectrograms lacked detailed information, leading to noticeable differences between the reconstructed vibrations and the originals.
We also incorporated GANs into our framework. Although GANs could generate spectrograms with more detail compared to autoencoders, they were prone to mode collapse, producing similar results for different inputs. This issue rendered GANs unsuitable for our optimization approach based on input vectors. Additionally, GANs generate results from random noise, resulting in a mixed latent space. As a result, the generated vibrations underwent drastic changes when users moved the slider continuously. Unlike the original DSS research, which focused on visual stimuli, our work emphasizes tactile tasks, which are less sensitive than visual stimuli. Consequently, these abrupt changes made it difficult for users to touch and compare vibrations, complicating the task of selecting an appropriate slider position.
We also compared generative models, CGAN and ACGAN, with conditional information support. Under our training framework, CGAN tended to produce extremely sparse output results or exhibited a tendency not to converge. ACGAN performed relatively better among the control group models we experimented with; however, it faced challenges when applied in the optimization process such as DSS. ACGAN can differentiate features between different classes and control them, but the uneven input vector format of label + latent vector is not conducive to the optimization process of DSS.
These problems make it difficult to identify models with similar performance for comparison with our TEXasGAN. Therefore, we have not provided the performance results or user studies of the control group in this regard.
@@ V-B Problems in Other Alternatives
V-C User Study
Some generated samples of Stage I are shown in Fig. 8 to show the generation performance.
In the figure, the leftmost column shows five categories of target vibrations, while the three columns on the right show the vibrations generated by three users for these targets. We selected one vibration from each category for demonstration. The spectrograms reveal that our TEXasGAN successfully produces vibration patterns similar to the target for certain categories to a certain extent. While some results are less ideal, the tendency to generate more uniform vibration patterns shows the consistency of the system. This consistency is a strength, although it can make replicating vibrations with significant temporal fluctuations, like Grass Fiber, more challenging. Moreover, the GAN structure greatly aids in generation, but some detail loss brought by the autoencoder-like structure still results in smoother outputs compared to real vibrations. Despite this, the system effectively replicates the general vibration characteristics to a certain extent, even if the intensity is slightly reduced compared to the original samples, such as Rough Paper.
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focus on analyzing the results of Stage II, the classification task. From Fig. 10 we can see an overview of the result of the classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations (experimental group, experiment 2), compared to about 72.8% for the real vibrations (control group, experiment 1). The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 10 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 kinds of different vibrations. This preliminary evidence suggests that our TEXasGAN can generate distinct vibration patterns that subjects can differentiate.
Fig. 11 shows the confusion matrix of the classification task of generated vibrations and Fig. 12 shows the control group, the classification of real vibrations. The confusion matrices provide a clearer view of how subjects distinguished between each type of vibration. We can see that both 2 confusion matrices show a diagonal trend.
Figure 13: The average accuracy of each subject.
Observing the confusion matrices, we find that their trends are similar but there are also some differences in accuracy appearing in categories G4 (26%), G6 (23%) and G8 (32%). In the experimental group, subjects classified the generated G4 samples with an accuracy of 53%, G6 samples with an accuracy of 46% and G8 samples with an accuracy of 37%, while in the control group, the accuracy for G6 and G8 was 79%, 69% and 69%, respectively. Actually, we can find that the reason is that more G4, G6, and G8 samples were classified as G2. According to subject feedback, the characteristics of G2 vibrations are less distinct compared to other categories. Subjects often described G2 as a uniform and smooth mid-low frequency vibration or found it difficult to describe any prominent patterns. We can also notice that even in the control group, distinguishing the real G2 vibrations was challenging, with an accuracy of 47%, same as the generated group.
We constructed a linear regression model with interaction terms, using control group accuracy, subject ID, and vibration class as independent variables, and the generated group accuracy as the response. Before the fit process, we removed outliers that were more than two scaled median absolute deviations from the median, obtained 93 sets of accuracy. The regressed model can be expressed as:
@@ V-C User Study
In the equation, $acc_{G}$ represents the generated group accuracy, $c$ represents the class ID, $i$ represents the subject ID and $acc_{R}$ represents the control group accuracy. The $p$-values of each term are shown in Table I. The analysis reveals that the accuracy of the control group, $acc_{R}$, has the most significant impact, along with the interaction term between class $c$ and $acc_{R}$, both of which are statistically significant. This suggests a meaningful relationship between $acc_{G}$ and $acc_{R}$, indicating that our model successfully reconstructs recognizable features of real target samples. The significance of the interaction term $c*acc_{R}$ implies that the model’s generation capability varies across different classes. Additionally, terms involving the subject ID $i$ are not significant, suggesting that the observed patterns are consistent across subjects.
TABLE I: Coefficients and $p$-values of each factor in the model.
@@
Then we performed a stepwise regression to select some variables. We can obtain a model like:
@@
This model keeps class $c$ ($p=0.0699$), control group accuracy $acc_{R}$ ($p=1.1785\times 10^{-8}$), and their interaction term $c*acc_{R}$ ($p=0.0012$). The $acc_{R}$ is still the most dominant term. The model still shows the difference in generation ability between classes, even larger than the previous model. This is reasonable according to our experiment results. The model $R^{2}=0.423$, means there is some certain correlation; $F=21.7$ with $p=1.19\times 10^{-10}$, means that the model is credited.
@@
Furthermore, we can also obtain some information from the classification in a class. The overall 3-class classification accuracy is 42.2%. This in-class accuracy of each class is shown in Fig. 14. The accuracy of each class is 45.6% (G2), 46.1% (G3), 41.1% (G4), 36.1% (G6) and 42.2% (G8). Although the accuracy for each class is higher than the random classification for three classes (33.3%), indicating that there are indeed some differences between samples generated with different targets, the classification accuracy is not high enough to prove that users can distinguish between samples belonging to the same class. This can be caused by the performance limitation of the model on the reconstruction of detailed information or the limited haptic resolution. However, in another aspect, this result suggests that the model captures and replicates some essential and general characteristics that define each class. Despite challenges in distinguishing within a class, participants were able to distinguish between different classes. This suggests that the system reliably encodes and reproduces the defining features of each kind of vibration.
@@
VI Discussion
From the results, we also observed several key phenomena that merit further discussion and analysis.
First, we can do some further discussion based on subjects’ feedback and the patterns of these types of vibrations. The main frequency component of G4 vibrations is similar to G3 while the amplitude is stronger, similar to G2. Although the accuracy of real G4 achieves a high level, many subjects said G4 is difficult to distinguish comparing to G2. Just some subjects paying attention to frequencies can point out that G4 is “smooth”, like G3 but less than G3. This observation helps explain why G2 and G4 are challenging to differentiate in the generated group.
This also suggests that subjects tend to use the intensity of the vibrations as their primary or initial criterion for the judgment. G6 vibrations are primarily composed of low-frequency components with uneven, strong, sudden pulses. And G8 vibrations are composed of some different frequency components, also strong but not so uneven as G6. Subjects who can pay attention to the frequency component can easily point out the differences but a lot of subjects just judge from the intensity and say they are similar. This phenomenon appears in both groups. Observing the spectrograms generated by users in Fig. 8 we notice that the model tends to produce uniform vibration patterns, making the generated vibrations similar to both G2 and G8. This similarity leads to confusion during classification as well.
Additionally, although the model incorporates GAN to mitigate detail loss due to feature space compression, some distortion is inevitable. As a result, the generated vibrations can exhibit an “oversmoothing” effect similar to that seen in autoencoders, making them less “sharp” compared to the real vibrations. This can also reduce the uneven pulses present in the vibrations. For the same reasons, the intensity of the originally strong and rough surface, G8, is weakened. This makes it more similar to the more uniform and weaker G2. This aligns with subject feedback and the observation mentioned earlier that subjects tend to prioritize amplitude when making initial judgments.
Our demonstrated model can generate tactile vibrations that are distinguishable and aligned with target characteristics according to the user’s preference. However, there is still an obvious gap between the classification task accuracy of real samples and generated samples. Although there are some differences between the vibration feedback and textures, vibrations can activate some related receptors and give some similar perceptions [6]. Therefore, some related conclusions about texture perceptions may can also explain phenomenons in this research. A previous study points out that, eliminating movement constricted the subjective range of texture roughness or other surface features [32]. Because the haptic vibrations presented in this study are time-related rather than space-related, the participants did not experience spatial movement during their perception process. However, vibrational and spatial stimulation is considered 2 components of texture perception [33]. The accuracy of classification task may be influenced by this kind of stimulation absence, especially in experiment 3 of the Stage II which focuses on the detailed differences in the same class.
Additionally, the phenomenon in which users can distinguish between different classes but struggle to differentiate samples within the same class might be due to the model capturing more general information while having difficulty in capturing more specific information. This may be another kind of mode collapse. The mode collapse can occur in one of two forms: the intra-class mode collapse and the inter-class mode collapse [34]. Although we have incorporated conditional information to enable the model to generate samples across multiple categories and to control the generation process, thereby avoiding inter-category mode collapse, intra-category mode collapse can still occur. This can lead to the condition occurring in Stage I, where subjects, while exploring similar feature spaces, encounter the model’s tendency to produce similar results for a specific class or neighboring features. Consequently, this results in the model’s inability to capture specific details accurately, leading to a lower distinction accuracy among subjects in experiment 3, Stage II.
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models could help improve the variety and fidelity of the output [35]. Diffusion models can be easily to train compared with GAN and can provide high quality results. Additionally, the current optimization process is quite time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation could significantly improve user experience. To further create realistic interaction experiences, we will consider using devices that better simulate texture perception, integrating both spatial and temporal stimulation.
@@ VII Conclusion
In conclusion, our study demonstrates the feasibility and effectiveness of a human-in-the-loop vibration generation system, TEXasGAN, to create realistic and distinguishable haptic textures. In our TEXasGAN model, we incorporate the idea of SRGAN and CAAE to generate class-controllable haptic vibrations with conditional information and encoded features. Using DSS and the pre-trained generative model, we enable users to intuitively control high-dimensional latent space and generate vibration samples using simple, one-dimensional sliders. The user experiments conducted with this system indicate that it can produce vibration patterns that are generally recognizable and distinguishable by subjects, though challenges remain in capturing finer, specific characteristics within specific classes. Overall, our approach offers a novel method for generating haptic vibrations based on user preferences, providing a foundation for future advancements in the field of haptic technology and virtual reality interactions.

### v4

@@ V-C User Study
In the figure, the leftmost column shows five categories of target vibrations, while the three columns on the right show the vibrations generated by three users for these targets. We selected one vibration from each category for demonstration. The spectrograms reveal that our TexSenseGAN successfully produces vibration patterns similar to the target for certain categories. While some results are less ideal, the tendency to generate more uniform vibration patterns shows the consistency of the system. This consistency is a strength, although it can make replicating vibrations with significant temporal fluctuations, such as Grass Fiber, more challenging. Moreover, the GAN structure significantly aids in generation; however, some detail loss owing to the autoencoder-like structure results in smoother outputs compared to real vibrations. Nonetheless, the system effectively replicates the general vibration characteristics to a certain extent, even if the intensity is slightly reduced compared to the original samples, such as Rough Paper.
However, spectrograms alone cannot fully capture the subjective experiences of the subjects. Therefore, we mainly focused on analyzing the results of Stage II, the classification task. From Fig. 12, we can observe an overview of the result of the classification task. In the five-category task, subjects achieved an accuracy of 55.9% when judging the generated vibrations (experimental group, experiment 2), compared to approximately 72.8% for the real vibrations (control group, experiment 1). The control group results indicate that subjects can distinguish between different surface vibration stimuli. Fig. 12 shows that subjects in the generation group and the control group exhibited similar accuracy trends for these 5 types of different vibrations. This preliminary evidence suggests that our TexSenseGAN can generate distinct vibration patterns that subjects can differentiate.
Fig. 13 shows the confusion matrix of the classification task of generated vibrations and Fig. 14 shows the control group, the classification of real vibrations. The confusion matrices provide a clearer view of how subjects distinguished between each type of vibration. We can observe that both 2 confusion matrices show a diagonal trend.
Observing the confusion matrices, we find that their trends are similar but there are some differences in accuracy appearing in categories G4 (26%), G6 (23%), and G8 (32%). In the experimental group, subjects classified the generated G4 samples with an accuracy of 53%, G6 samples with an accuracy of 46% and G8 samples with an accuracy of 37%, while in the control group, the accuracy for G4, G6, and G8 was 79%, 69%, and 69%, respectively. This is because more G4, G6, and G8 samples were classified as G2. According to subject feedback, the characteristics of G2 vibrations are less distinct compared to other categories. Subjects often described G2 as a uniform and smooth mid-low frequency vibration or found it difficult to describe any prominent patterns. It can be observed that even in the control group, distinguishing the real G2 vibrations was challenging, with an accuracy of 47%, same as the generated group.
@@
We constructed a linear regression model with interaction terms, using control group accuracy, subject ID, and vibration class as independent variables, and the generated group accuracy as the response. Before the fit process, we removed outliers that exceeded two scaled median absolute deviations from the median, obtained 93 sets of accuracy. The regressed model can be expressed as:
@@
In this equation, $acc_{G}$ represents the generated group accuracy, $c$ represents the class ID, $i$ represents the subject ID, and $acc_{R}$ represents the control group accuracy. The $p$-values of each term are shown in Table II. The analysis reveals that the accuracy of the control group, $acc_{R}$, has the most significant impact, along with the interaction term between class $c$ and $acc_{R}$, both of which are statistically significant. This suggests a meaningful relationship between $acc_{G}$ and $acc_{R}$, indicating that our model successfully reconstructs recognizable features of real target samples. The significance of the interaction term $c*acc_{R}$ implies that the model’s generation capability varies across different classes. Additionally, terms involving the subject ID $i$ are insignificant, suggesting that the observed patterns are consistent across subjects.
Figure 15: The average accuracy of each subject.
TABLE II: Coefficients and $p$-values of each factor in the model.
@@
We performed a stepwise regression to select some variables. We can obtain a model like:
@@
This model maintains class $c$ ($p=0.0699$), control group accuracy $acc_{R}$ ($p=1.1785\times 10^{-8}$), and their interaction term $c*acc_{R}$ ($p=0.0012$). The $acc_{R}$ is the most dominant term. The model shows the difference in generation ability between classes, even larger than the previous model. This is reasonable according to our experiment results. The model $R^{2}=0.423$, reveals that there is a correlation; $F=21.7$ with $p=1.19\times 10^{-10}$, indicating that the model is credited.
@@
Furthermore, we can obtain some information from the classification in a class. The overall 3-class classification accuracy is 42.2%. This in-class accuracy of each class is shown in Fig. 16. The accuracy of each class is 45.6% (G2), 46.1% (G3), 41.1% (G4), 36.1% (G6) and 42.2% (G8). Although the accuracy for each class is higher than the random classification for three classes (33.3%), indicating that there are indeed some differences between samples generated with different targets, the classification accuracy is not sufficiently high to prove that users can distinguish between samples belonging to the same class. This can be caused by the performance limitation of the model on the reconstruction of detailed information or the limited haptic resolution. However, in another aspect, this result suggests that the model captures and replicates some essential and general characteristics that define each class. Despite challenges in distinguishing within a class, participants could distinguish between different classes. Therefore, the system reliably encodes and reproduces the defining features of each type of vibration.
VI Discussion
From the results, we observed several key phenomena that merit further discussion and analysis.
First, we can further discuss based on subjects’ feedback and the patterns of these types of vibrations. The main frequency component of G4 vibrations is similar to G3 while the amplitude is stronger, similar to G2. Although the accuracy of real G4 achieves a high level, various subjects said G4 is difficult to distinguish comparing to G2. Just some subjects paying attention to frequencies can point out that G4 is “smooth”, such as G3 but less than G3. This observation helps explain why G2 and G4 are challenging to differentiate in the generated group.
This also suggests that subjects tend to use the intensity of the vibrations as their primary or initial criterion for the judgment. G6 vibrations are typically composed of low-frequency components with uneven, strong, sudden pulses. G8 vibrations are composed of some different frequency components, also strong but not uneven as G6. Subjects who can pay attention to the frequency component can easily point out the differences but many subjects just judge from the intensity and say they are similar. This phenomenon appears in both groups. Observing the waveforms generated by users in Fig. 11 we notice that the model tends to produce uniform vibration patterns, making the generated vibrations similar to both G2 and G8. This similarity causes confusion during classification as well.
Additionally, although the model incorporates GAN to mitigate detail loss owing to feature space compression, some distortion is inevitable. Consequently, the generated vibrations can exhibit an “oversmoothing” effect similar to that observed in autoencoders, making them less “sharp” compared to the real vibrations. This can reduce the uneven pulses present in the vibrations. Similarly, the intensity of the originally strong and rough surface, G8, is weakened. Thus, it is more similar to the more uniform and weaker G2. This aligns with subject feedback and the observation mentioned earlier that subjects tend to prioritize amplitude when making initial judgments.
Our demonstrated model can generate tactile vibrations that are distinguishable and aligned with target characteristics according to the user’s preference. However, there is a gap between the classification task accuracy of real samples and generated samples. Although there are some differences between the vibration feedback and textures, vibrations can activate some related receptors and provide some similar perceptions [7]. Therefore, some related conclusions about texture perceptions may explain phenomenons in this research. A previous study points out that, eliminating movement constricted the subjective range of texture roughness or other surface features [34]. Because the haptic vibrations presented in this study are time-related rather than space-related, the participants did not experience spatial movement during their perception process. However, vibrational and spatial stimulation considered two components of texture perception [35]. The accuracy of classification task may be influenced by this type of stimulation absence, especially in experiment 3 of Stage II, which focuses on the detailed differences in the same class.
Additionally, the phenomenon in which users can distinguish between different classes but struggle to differentiate samples within the same class might be owing to the model capturing more general information while having difficulty in capturing more specific information. This may be another type of mode collapse. The mode collapse can occur in one of two forms: the intra-class mode collapse and the inter-class mode collapse [36]. Although we have incorporated conditional information to enable the model to generate samples across multiple categories and to control the generation process, thereby avoiding inter-category mode collapse, intra-category mode collapse can occur. This can lead to the condition occurring in Stage I, where subjects, while exploring similar feature spaces, encounter the model’s tendency to produce similar results for a specific class or neighboring features. Consequently, this results in the model’s inability to capture specific details accurately, leading to a lower distinction accuracy among subjects in experiment 3, Stage II.
Therefore, in future work, we plan to explore other generative models to enhance the diversity and precision of the generated vibrations. For example, incorporating diffusion models can help improve the variety and fidelity of the output [37]. Diffusion models can be easy to train compared with GAN and can provide high quality results. Additionally, the current optimization process is time consuming and requires subjects to focus intensely on detecting subtle differences while adjusting the slider position. Streamlining the interface for more intuitive operation can significantly improve user experience. Our current presentation system is designed specifically to deliver texture-related vibratory feedback on simple devices, such as game controllers, rather than to replicate the sensation of real textures themselves. Therefore, in future research, we may further explore the application of this framework to spatial and temporal stimuli and aim to deploy it on equipment better suited to the original data. For instance, we envision using devices like the Phantom of 3D Systems, to fit the data collection method better and consider spatial stimuli to create more realistic tactile feedback.
@@ VII Conclusion
Our study demonstrates the feasibility and effectiveness of a human-in-the-loop vibration generation system, TexSenseGAN, to create realistic and distinguishable haptic textures. In our TexSenseGAN model, we incorporated the idea of SRGAN and CAAE to generate class-controllable haptic vibrations with conditional information and encoded features. By employing DSS and the pre-trained generative model, we enable users to intuitively control high-dimensional latent space and generate vibration samples using simple, one-dimensional sliders. The user experiments conducted with this system indicate that it can produce vibration patterns that are recognizable and distinguishable by subjects, though challenges remain in capturing finer, specific characteristics within specific classes. Our approach offers a novel method for generating haptic vibrations based on user preferences, providing a foundation for future advancements in the field of haptic technology and virtual reality interactions.